A man borrowed P200,000 from a bank at 12% compounded monthly, which is payable monthly for 10 years (120 payments). If the first payment is to be made after 3 months, how much is the monthly payment?



1
Expert's answer
2022-03-15T16:50:45-0400

The monthly payment is given as:

"Pmt=\frac{PV}{[\frac{1-\frac{1}{(1+\frac{r}{m})^{nm}}}{\frac{r}{m}}]}"

Where PV is P200000

r is 12%

n is 10

m is 12

"\therefore Pmt=\frac{200000}{[\frac{1-\frac{1}{(1+\frac{12\%}{12})^{10×12}}}{\frac{12\%}{12}}]}"

"Pmt=P286.94"